# Perceptions of the Fairness Impacts of Multiplicity in Machine Learning

### v1

@@ 3 Related Work and Hypotheses
3.1 Algorithm aversion
Algorithm aversion is when people prefer human decision-makers, even when the available AI system generally outperforms humans [27]. It may be triggered when people see a model make a mistake, even if they would have been forgiving of a human making the same mistake [27]. Algorithm aversion also tends to be higher for subjective tasks, a characterization that is malleable depending on how the task is described [15]. We expect that witnessing model multiplicity will increase algorithm aversion for three reasons: first, the existence of two good (but different) models means that the models make mistakes. Second, the existence of multiple good models indicates that the decisions these models make are subjective. Finally, existing work assumes (without consulting stakeholders) that multiplicity threatens fairness (section 2.2).
@@ 3.1 Algorithm aversion
Between algorithm aversion and prior work showing preference for human over AI arbitrators [86], we think participants will prefer human over other resolution techniques (recall that human and the other approaches are defined in table 1). We think that – counter to what philosophical work suggests – random will be viewed as the least fair approach. Explanations are important to fairness [83, 101], so we think the sense of unpredictability from randomized decisions dominate stakeholders’ reactions to random.
@@
3.2 Task stakes
Competing with algorithm aversion is algorithm appreciation, when people over-rely on algorithmic decisions [60]. There is evidence that algorithm appreciation occurs in high-stakes tasks [1, 81], and stakeholders rate accuracy as more important in high-stakes settings [90]. Thus, we think that fancy (i.e., building a more complex model) may be preferred for high-stakes tasks because it may be perceived as being more accurate. However, the importance of accuracy may also trigger algorithm aversion and increase preference for human. Fairness is more important to stakeholders in high-stakes settings than low-stake ones [45]. We suspect that higher-stakes tasks are more emotionally charged. People view emotional decision making as a strength of humans [39], and can have negative emotional responses to AI decision-making [56].
@@ 3.2 Task stakes
3.3 Task framing
Harm that is caused by action, rather than omission, causes harsher fairness judgments [24]. Tasks that are punitive commit active harm in the event of a mistake, whereas reward-based tasks cause harm when they make an omission. Thus, we think that punitive tasks will have greater decreases in fairness perceptions in the presence of multiplicity. Indeed, others have found that when people adopt critical mindsets, they are more likely to find issues with a model [37]. Algorithms are seen as more rational than humans [60], and we suspect that rationality will be more important in punitive settings where mistakes are judged more harshly. Conversely, there are positive connotations towards human accuracy [56], so being primed to think about reward settings may influence participants to select human.
@@ 3.3 Task framing
3.4 Task uncertainty
High-uncertainty tasks typically do not have a single objectively-right answer, such as predicting future human behavior (e.g., job performance) or matters of taste (e.g., movie recommendation). Low-uncertainty tasks typically have an objectively-correct answer (e.g., detecting spam emails). Algorithm appreciation is known to be stronger in objective (i.e., low-uncertainty) tasks [2, 60], aligning with beliefs that algorithms are not good at subjective (i.e., high-uncertainty) tasks [15]. There is more algorithm aversion and preference for human decision makers in high-uncertainty settings [26]. When a task has lower uncertainty, it may be possible to have overall higher accuracy, which increases algorithm appreciation [81]. Intuitively, people may find it easier to rationalize that there are multiple good solutions to a high-uncertainty task, resulting in lower impact to the perceived fairness.
@@ 3.4 Task uncertainty
4 Method
Figure 2 shows an overview of our study design, which was approved by our institution’s IRB board. In this section, we first describe our preliminary study (section 4.1), which informed the task selection for our main study. Then, we describe our pre-registered study design (section 4.2) and data analysis plan (section 4.4).
4.1 Task selection (Preliminary study)
@@
We collect 18 scenarios through brainstorming and borrowing from previous studies [6, 15, 48]. We randomly assign each participant to see 4 scenarios to limit participant fatigue when completing the study. For each scenario, participants answer a series of 5-point Likert-scale questions relating to task clarity (2 questions), task stakes (3 questions), and task uncertainty (2 questions). Then, participants rank the four tasks according to stakes and uncertainty levels. We use Prolific to find respondents (US-based adults with a 97%+ approval rate). We compensate participants at Prolific’s recommended rate ($12/hour) based on the estimated survey completion time.
@@
Our survey yields four pieces of data: stakes rating, uncertainty rating, stakes ranking, and uncertainty ranking. We convert all Likert ratings to a scale from -2 (strongly disagree) to 2 (strongly agree). For the rating data, we average the three questions relating to task stakes into a single stakes rating quantity. We elect to use one of the task uncertainty questions alone, due to better data quality. To recover the underlying stakes and uncertainty rankings given the partial ranking information from each participant, we use assume a Plackett-Luce distribution [62, 73] and use the I-LSR (iterative Luce Spectral Rating) algorithm [66]. The basic assumption behind Plackett-Luce is that choice rankings are transitive (i.e., if A¿B and B¿C then A¿C), and the I-LSR algorithm uncovers the full underlying ranking. We combine the rating and ranking data into a single metric by scaling each set of data to have mean 0 and variance 1, and then taking their arithmetic mean.
@@
4.2 Study design
The survey flow is depicted in Figure 2. The survey begins with education about multiplicity (with a comprehension question) and then has two main sections (“Part 1” and “Part 2”). At the end, we collect demographic information. Participants are randomly assigned to one task at the beginning of Part 1. Then we describe a scenario and solicit the participant’s fairness perceptions of the model on a Likert scale consisting of four seven-point questions. In Part 2, we present a series of binary choices between different ways to resolve multiplicity and the user selects their preferred option. The demographic information we collect includes age, gender, education level, and experience with machine learning (defined as whether the participant uses ML, now or in the past, as part of their work or school). We provide further details on all parts of the survey below.
Details of comprehension question
The comprehension question describes multiplicity using a fictional, low-stakes example with text and an image similar to Figure 1B. Then we ask participants to select the correct definition of multiplicity from a list of choices. The full text of the comprehension question is available in the appendix.
@@ 4.2 Study design
Next, we ask users to respond to four seven-point Likert scale questions to elicit fairness perceptions. Following other studies that ask about multiple dimensions of fairness [20, 56, 83, 101], we use two questions relating to distributional fairness and two relating to procedural fairness.
Then, we describe how multiplicity impacts the situation, explaining that there is a second model that agrees with human experts equally often as the first model, but disagrees less often with the original model. We accompany the description with a graphic such as the one shown in Figure 1C. At this point, participants answer two critical-thinking questions: one is multiple-choice and comprehension based; the other is open-ended and asks how decisions should be made when the two models disagree.
The next page of the survey asks the participant to respond again to the same set of Likert questions. We refer to the first set of responses as the “pre” timestamp, and this set of responses as the “post” timestamp, because the latter are collected after the participant learns that multiplicity exists for the task. The participant does not have access to their previous responses at the “post” timestamp, because we do not want them to be anchored by their original responses (i.e., we want to allow for a change in their frame of reference [43]).
@@
4.3 Data collection
We collect data using Prolific. Our participation criteria is US-based adults (18+) who have at least a 95% approval rate. We plan to collect data from 360 participants, i.e., 45 per task. We estimated that 35-40 participants per task would be sufficient based on trends in pilot data; we rounded up to 45/task to have additional buffer in case we need to reject some samples. We follow Prolific’s guidance [74] and exclude survey respondents who fail the initial comprehension check question twice, fail two or more attention check questions, or are statistical outliers in survey completion time (in the fast direction).
4.4 Statistical analyses
4.4.1 Part 1 analysis
@@
To test H-H, we fit a linear mixed model (LMM), which generalizes a linear model by accounting for both between- and within-subject factors. The formula we fit to the LMM predicts the transformed Likert data as a function of time, task stakes, task framing, task uncertainty, respondent ID (to account for individual differences), and the first-order interaction terms of time with each task characteristic. We use the lme4 package [4] in R to construct the LMM model and then use the car package [33] to conduct a type-III ANOVA. Using ANOVA procedures on Likert-scale data is controversial because Likert scales are not continuous and may not produce normally distributed data. Scholars have argued both for [9, 14, 70] and against [46, 54] applying ANOVA when its assumptions are violated. We decided that LMMs fit our needs best, so to minimize concerns about using ANOVA, we included a large population in each group (as required by the central limit theorem for normality of the data sample) and used the average of several Likert items rather than a single item [14]. The main alternative technique we considered is Aligned-Rank-Transform (ART) ANOVA [98], a procedure that transforms ordinal (e.g., Likert-scale) data to fit the ANOVA assumptions. Ultimately we chose LMMs because ART-ANOVA does not allow for continuous factors. However, for readers who are interested, we include two additional analyses in the appendix: ART-ANOVA and a non-parametric test, the Mann-Whitney U-Test [64].333To clarify, we preregistered the analysis with LMMs. Readers should view the results of the other tests as a partial multiverse analysis [92].
To test H, we check whether the ANOVA p-value is significant, and if so, run a post-hoc contrast test to gauge directionality. To test H and H, we find the estimated marginal trends for how task stakes (H) and uncertainty (H) vary with time using emtrends in the emmeans package in R [85]. Then, we conduct a pairwise contrast test on the result. To test H, we perform a pairwise contrast test on the estimated marginal means of the interaction term of time and task framing, again using emmeans.
@@
In total, since we perform four post-hoc contrast tests on the Likert scale data, we will use the Holm-Bonferroni correction [42] with and to evaluate the hypotheses.
4.4.2 Part 2 analysis
@@
We test H-H using choice-based conjoint analysis [36]. Choice-based analysis is a technique that solicits user preferences over two options with various factors set to different attribute levels. The analysis generates the marginal means of each attribute level. Our analysis uses the cregg package in R [58]. For H, we look directly at the marginal mean effect of each of the multiplicity resolution techniques and costs. For each condition, the marginal mean procedure outputs a point estimate and a confidence interval of how preferred an option is relative to random chance. For instance, an output of 0.8 indicates that participants presented with that option chose it 80% of the time. For H, H, and H, we compute the marginal means stratified by task characteristics. To the best of our knowledge, there is no tool to compute marginal means for a continuous factor, so we use the binary encodings (low vs. high) for task stakes and task uncertainty.
@@
We will consider H to be weakly supported if the confidence interval for human is strictly greater than 0.5 and the confidence interval for random is strictly less than 0.5. Then, we will test whether the estimate for human is strictly higher than the estimates for the other methods. To do this, we will perform five two-sided t-tests444The choice-based analysis data may not be normally distributed; however, we have over 800 observations for each resolution technique so the central limit theorem makes the use of this test acceptable. (one against each other multiplicity resolution technique). Then, we will perform four t-tests (random vs. each other multiplicity resolution technique other than human) to see whether random is strictly less preferred than each other technique.
@@
Since there are six multiplicity resolution techniques, we use the Bonferroni555Note, we do not use Holm-Bonferroni here because it is not supported by cregg. correction to get a p-value of and use this to compute the confidence intervals for each marginal means estimate. We perform a total of 18 post-hoc t-tests on the data, so will evaluate their significance using the Holm-Bonferroni correction with and .
5 Data Analysis
5.1 Summary of responses

### v2

@@ 3. Related Work and Hypotheses
3.1. Algorithm aversion
Algorithm aversion is when people prefer human decision-makers, even when the available AI system generally outperforms humans (Dietvorst et al., 2015). It may be triggered when people see a model make a mistake, even if they would have been forgiving of a human making the same mistake (Dietvorst et al., 2015), or when the model’s performance is worse than stated (Yin et al., 2019). In settings where a correct answer can take many forms (e.g., generative AI), seeing multiple conflicting output lowers trust in a model (Lee et al., 2024). Algorithm aversion also tends to be higher for subjective tasks, a characterization that is malleable depending on how the task is described (Castelo et al., 2019). We expect that witnessing model multiplicity will increase algorithm aversion for three reasons: first, the existence of two good (but different) models means that the models make mistakes, similar to what was observed for conflicting generative AI responses (Lee et al., 2024). Second, the existence of multiple good models indicates that the decisions these models make are subjective. In particular, multiplicity means that choosing a model is not a purely objective task (i.e., optimizing accuracy), but rather a subjective choice of choosing between multiple models with near-equivalent accuracy. Finally, existing work assumes (without consulting stakeholders) that multiplicity threatens fairness (section 2.2).
@@ 3.1. Algorithm aversion
Between algorithm aversion and prior work showing preference for human over AI arbitrators (Sela, 2018), we think participants will prefer human over other resolution techniques (recall that human and the other approaches are defined in table 1). We think that – counter to what philosophical work suggests – random will be viewed as the least fair approach. Explanations are important to fairness (Schoeffer et al., 2022; Yurrita et al., 2023), so we think the sense of unpredictability from randomized decisions dominate stakeholders’ reactions to random.
@@
3.2. Task stakes
Competing with algorithm aversion is algorithm appreciation, when people over-rely on algorithmic decisions (Logg et al., 2019). There is evidence that algorithm appreciation occurs in high-stakes tasks (Araujo et al., 2020; Saragih and Morrison, 2022), and stakeholders rate accuracy as more important in high-stakes settings (Srivastava et al., 2019). Thus, we think that fancy (i.e., building a more complex model) may be preferred for high-stakes tasks because it may be perceived as being more accurate. However, the importance of accuracy may also trigger algorithm aversion and increase preference for human. Fairness is more important to stakeholders in high-stakes settings than low-stake ones (Jakesch et al., 2022). We suspect that higher-stakes tasks are more emotionally charged. People view emotional decision making as a strength of humans (Helberger et al., 2020), and can have negative emotional responses to AI decision-making (Lee, 2018).
@@ 3.2. Task stakes
3.3. Task framing
Harm that is caused by action, rather than omission, causes harsher fairness judgments (Cushman et al., 2006). Tasks that are punitive commit active harm in the event of a mistake, whereas reward-based tasks cause harm when they make an omission. Thus, we think that punitive tasks will have greater decreases in fairness perceptions in the presence of multiplicity. Indeed, others have found that when people adopt critical mindsets, they are more likely to find issues with a model (Hannan et al., 2021). Algorithms are seen as more rational than humans (Logg et al., 2019), and we suspect that rationality will be more important in punitive settings where mistakes are judged more harshly. Conversely, there are positive connotations towards human accuracy (Lee, 2018), so being primed to think about reward settings may influence participants to select human.
@@ 3.3. Task framing
3.4. Task uncertainty
High-uncertainty tasks typically do not have a single objectively-right answer, such as predicting future human behavior (e.g., job performance) or matters of taste (e.g., movie recommendation). Low-uncertainty tasks typically have an objectively-correct answer (e.g., detecting spam emails). Algorithm appreciation is known to be stronger in objective (i.e., low-uncertainty) tasks (Bankins et al., 2022; Logg et al., 2019), aligning with beliefs that algorithms are not good at subjective (i.e., high-uncertainty) tasks (Castelo et al., 2019). There is more algorithm aversion and preference for human decision makers in high-uncertainty settings (Dietvorst and Bharti, 2020). When a task has lower uncertainty, it may be possible to have overall higher accuracy, which increases algorithm appreciation (Saragih and Morrison, 2022). Intuitively, people may find it easier to rationalize that there are multiple good solutions to a high-uncertainty task, resulting in lower impact to the perceived fairness.
@@ 3.4. Task uncertainty
4. Method
Figure 2 shows an overview of our study design, which was approved by our institution’s IRB board. In this section, we first describe our preliminary study (section 4.1), which informed the task selection for our main study. Then, we describe our pre-registered study333https://osf.io/fv3bt design (section 4.2) and data analysis plan (section 4.4).
Visual overview of our study design. There are two rows. First row text is Preliminary study (n=199). Goal: determine task stakes and uncertainty to select tasks for the main study. There are 18 tasks total and each person sees 4. We ask How high-stakes is the task (Likert), How much uncertainty is there in the task outcomes (Likert), and we ask the participants to rank the 4 tasks they see on stakes and on uncertainty. The second row says Main Study (n=358). Goal: determine how multiplicity impacts fairness (Part1) and how participants think multiplicity should be resolved (Part 2). There are 8 tasks total and each person sees 1. The remainder of this row is subdividing into Part 1 and Part 2. In part 1, the flow is Comprehension (education on what multiplicity is) to Initial fairness perceptions (Likert) to Critical thinking (reveal that multiplicity exists, open-ended response), to Final fairness perceptions (Likert). In Part 2, it says Choice Based Analysis (we ask do you prefer option A or Option B). We randomly populate a table with two rows and two columns. The first column is labelled option A and the second column is labelled option B. The first row is How to resolve multiplicity and the second row is Cost. The example shown in the image has Option A= Choose randomly and None and Option B = Human employee and 5 dollars.
4.1. Task selection (Preliminary study)
@@
We collect 18 scenarios through brainstorming and borrowing from previous studies (Binns et al., 2018; Castelo et al., 2019; Kasinidou et al., 2021). We randomly assign each participant to see 4 scenarios to limit participant fatigue when completing the study. For each scenario, participants answer a series of 5-point Likert-scale questions relating to task clarity (2 questions), task stakes (3 questions), and task uncertainty (2 questions). Then, participants rank the four tasks according to stakes and uncertainty levels. We use Prolific to find respondents (US-based adults with a 97%+ approval rate). We compensate participants at Prolific’s recommended rate ($12/hour) based on the estimated survey completion time.
@@
Our survey yields four pieces of data: stakes rating, uncertainty rating, stakes ranking, and uncertainty ranking. We convert all Likert ratings to a scale from -2 (strongly disagree) to 2 (strongly agree). For the rating data, we average the three questions relating to task stakes into a single stakes rating quantity. We elect to use one of the task uncertainty questions alone, due to better data quality. To recover the underlying stakes and uncertainty rankings given the partial ranking information from each participant, we use assume a Plackett-Luce distribution (Luce, 1977; Plackett, 1975) and use the I-LSR (iterative Luce Spectral Rating) algorithm (Maystre and Grossglauser, 2015). The basic assumption behind Plackett-Luce is that choice rankings are transitive (i.e., if A¿B and B¿C then A¿C), and the I-LSR algorithm uncovers the full underlying ranking. We combine the rating and ranking data into a single metric by scaling each set of data to have mean 0 and variance 1, and then taking their arithmetic mean.
@@
4.2. Study design
The survey flow is depicted in Figure 2. The survey begins with education about multiplicity (with a comprehension question) and then has two main sections (“Part 1” and “Part 2”). Participants are randomly assigned to one task at the beginning of Part 1. Then we describe a scenario and solicit the participant’s fairness perceptions of the model on a Likert scale consisting of four seven-point questions. In Part 2, we present a series of binary choices between different ways to resolve multiplicity and the user selects their preferred option. We conclude by collecting demographic information from the participant. We provide further details on all parts of the survey below.
Details of multiplicity education
We describe to participants, at a high level, how AI models are built (e.g., “AI developers […] choose the model with the best performance”). Then, we reveal that there are sometimes multiple models that can perform equally well, but “make different mistakes”. We illustrate model multiplicity using a graphic similar to fig. 1B. We tell participants that this phenomenon, model multiplicity, is common. Then, participants have two chances to correctly answer the multiple-choice question “What does model multiplicity mean in the context of AI?” The correct answer is written using similar phrasing as in the descriptive text. We include full text of the multiplicity education section in the appendix.
@@ 4.2. Study design
Next, we ask users to respond to four seven-point Likert scale questions to elicit fairness perceptions. Two questions focus on the model outcomes, e.g., whether the proposed model is likely to be able to make “good” decisions, and two questions focus on whether using the model is appropriate. Then, we describe how multiplicity impacts the situation, explaining that there is a second model that agrees with human experts equally often as the first model, but disagrees less often with the original model. We accompany the description with a graphic such as the one shown in Figure 1C. At this point, participants answer two critical-thinking questions: one is multiple-choice and comprehension based; the other is open-ended and asks how decisions should be made when the two models disagree.
The next page of the survey asks the participant to respond again to the same set of Likert questions. We refer to the first set of responses as the “pre” timestamp, and this set of responses as the “post” timestamp, because the latter are collected after the participant learns that multiplicity exists for the task. The participant does not have access to their previous responses at the “post” timestamp, because we do not want them to be anchored by their original responses (i.e., we want to allow for a change in their frame of reference (Howard et al., 1979)).
@@
Demographic information
We collected demographic information from participants including gender, age, education level, and experience with AI. Experience with AI was collected through the question “Do you interact with artificial intelligence or automated decision-making systems as part of your work or school?”, with the choices “yes,” “no,” and “not currently, but in the past”.
4.3. Data collection
We collect data using Prolific. Our participation criteria is US-based adults (18+) who have at least a 95% approval rate. We plan to collect data from 360 participants, i.e., 45 per task. We estimated that 35-40 participants per task would be sufficient based on trends in pilot data; we rounded up to 45/task to have additional buffer in case we need to reject some samples. We follow Prolific’s guidance (Prolific, 2024) and exclude survey respondents who fail the initial comprehension check question twice, fail two or more attention check questions, or are statistical outliers in survey completion time (in the fast direction).
4.4. Statistical analyses
4.4.1. Part 1 analysis
@@
To test H-H, we fit a linear mixed model (LMM), which generalizes a linear model by accounting for both between- and within-subject factors. The formula we fit to the LMM predicts the transformed Likert data as a function of time, task stakes, task framing, task uncertainty, respondent ID (to account for individual differences), and the first-order interaction terms of time with each task characteristic. We use the lme4 package (Bates et al., 2015) in R to construct the LMM model and then use the car package (Fox and Weisberg, 2019) to conduct a type-III ANOVA. Using ANOVA procedures on Likert-scale data is controversial because Likert scales are not continuous and may not produce normally distributed data. Scholars have argued both for (Blanca Mena et al., 2017; Carifio and Perla, 2008; Norman, 2010) and against (Jamieson, 2004; Kuzon et al., 1996) applying ANOVA when its assumptions are violated. We decided that LMMs fit our needs best, so to minimize concerns about using ANOVA, we included a large population in each group (as required by the central limit theorem for normality of the data sample) and used the average of several Likert items rather than a single item (Carifio and Perla, 2008). The main alternative technique we considered is Aligned-Rank-Transform (ART) ANOVA (Wobbrock et al., 2011), a procedure that transforms ordinal (e.g., Likert-scale) data to fit the ANOVA assumptions. Ultimately we chose LMMs because ART-ANOVA does not allow for continuous factors. However, for readers who are interested, we include two additional analyses in the appendix: ART-ANOVA and a non-parametric test, the Mann-Whitney U-Test (Mann and Whitney, 1947).444To clarify, we preregistered the analysis with LMMs. Readers should view the results of the other tests as a partial multiverse analysis (Steegen et al., 2016).
To test H, we check whether the ANOVA p-value is significant, and if so, run a post-hoc contrast test to gauge directionality. To test H and H, we find the estimated marginal trends for how task stakes (H) and uncertainty (H) vary with time using emtrends in the emmeans package in R (Searle et al., 1980). Then, we conduct a pairwise contrast test on the result. To test H, we perform a pairwise contrast test on the estimated marginal means of the interaction term of time and task framing, again using emmeans.
@@
In total, since we perform four post-hoc contrast tests on the Likert scale data, we will use the Holm-Bonferroni correction (Holm, 1979) with and to evaluate the hypotheses.
4.4.2. Part 2 analysis
@@
We test H-H using choice-based conjoint analysis (Hainmueller et al., 2014). Choice-based analysis is a technique that solicits user preferences over two options with various factors set to different attribute levels. The analysis generates the marginal means of each attribute level. Our analysis uses the cregg package in R (Leeper, 2020). For H, we look directly at the marginal mean effect of each of the multiplicity resolution techniques and costs. For each condition, the marginal mean procedure outputs a point estimate and a confidence interval of how preferred an option is relative to random chance. For instance, an output of 0.8 indicates that participants presented with that option chose it 80% of the time. For H, H, and H, we compute the marginal means stratified by task characteristics. To the best of our knowledge, there is no tool to compute marginal means for a continuous factor, so we use the binary encodings (low vs. high) for task stakes and task uncertainty.
@@
We will consider H to be weakly supported if the confidence interval for human is strictly greater than 0.5 and the confidence interval for random is strictly less than 0.5. Then, we will test whether the estimate for human is strictly higher than the estimates for the other methods. To do this, we will perform five two-sided t-tests555The choice-based analysis data may not be normally distributed; however, we have over 800 observations for each resolution technique so the central limit theorem makes the use of this test acceptable. (one against each other multiplicity resolution technique). Then, we will perform four t-tests (random vs. each other multiplicity resolution technique other than human) to see whether random is strictly less preferred than each other technique.
@@
Since there are six multiplicity resolution techniques, we use the Bonferroni666Note, we do not use Holm-Bonferroni here because it is not supported by cregg. correction to get a p-value of and use this to compute the confidence intervals for each marginal means estimate. We perform a total of 18 post-hoc t-tests on the data, so will evaluate their significance using the Holm-Bonferroni correction with and .
5. Data Analysis